Explain how the rate of diffusion depends on the density of a gas


1
Expert's answer
2021-04-26T17:08:34-0400

According to Graham's law of diffusion, the following is true for the rates of diffusion "R" of two gases of different molecular weight "\Mu":

"\frac{R_1}{R_2}=\sqrt{\frac{M_2}{M_2}}."

As we see, the rate of diffusion of a gas is inversely proportional to the square root of its molecular mass. Taking equal volumes of gases, we see that the rate of diffusion is inversely proportional to the square root of the mass density of this gas.